OpenStudy (anonymous):

Find the product. y^5 · y^3 y =

OpenStudy (anonymous):

is it 8 or 15?

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (shubhamsrg):

a^n . a^m = a^(n+m)

OpenStudy (anonymous):

oh so its y^8?

OpenStudy (kendricklamar2014):

yes

OpenStudy (anonymous):

ok thnx:)

OpenStudy (kendricklamar2014):

y*y = y 5+3 = 8

OpenStudy (kendricklamar2014):

No problem